\begin{thebibliography}{36}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Aharoni and Goldberg(2020)}]{aharoni2020unsupervised}
Roee Aharoni and Yoav Goldberg. 2020.
\newblock \href {https://arxiv.org/abs/2004.02105} {Unsupervised domain
  clusters in pretrained language models}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}. Association for
  Computational Linguistics.

\bibitem[{Barbu et~al.(2016)Barbu, Parra~Escart\'{\i}n, Bentivogli, Negri,
  Turchi, Orasan, and Federico}]{10.1007/s10590-016-9183-x}
Eduard Barbu, Carla Parra~Escart\'{\i}n, Luisa Bentivogli, Matteo Negri, Marco
  Turchi, Constantin Orasan, and Marcello Federico. 2016.
\newblock \href {https://doi.org/10.1007/s10590-016-9183-x} {The first
  automatic translation memory cleaning shared task}.
\newblock \emph{Machine Translation}, 30(3–4):145–166.

\bibitem[{Belinkov and Bisk(2018)}]{DBLP:conf/iclr/BelinkovB18}
Yonatan Belinkov and Yonatan Bisk. 2018.
\newblock \href {https://openreview.net/forum?id=BJ8vJebC-} {Synthetic and
  natural noise both break neural machine translation}.
\newblock In \emph{6th International Conference on Learning Representations,
  {ICLR} 2018, Vancouver, BC, Canada, April 30 - May 3, 2018, Conference Track
  Proceedings}. OpenReview.net.

\bibitem[{Bulte and Tezcan(2019)}]{bulte-tezcan-2019-neural}
Bram Bulte and Arda Tezcan. 2019.
\newblock \href {https://doi.org/10.18653/v1/P19-1175} {Neural fuzzy repair:
  Integrating fuzzy matches into neural machine translation}.
\newblock In \emph{Proceedings of the 57th Annual Meeting of the Association
  for Computational Linguistics}, pages 1800--1809, Florence, Italy.
  Association for Computational Linguistics.

\bibitem[{Cai et~al.(2021)Cai, Wang, Li, Lam, and Liu}]{cai-etal-2021-neural}
Deng Cai, Yan Wang, Huayang Li, Wai Lam, and Lemao Liu. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.acl-long.567} {Neural machine
  translation with monolingual translation memory}.
\newblock In \emph{Proceedings of the 59th Annual Meeting of the Association
  for Computational Linguistics and the 11th International Joint Conference on
  Natural Language Processing (Volume 1: Long Papers)}, pages 7307--7318,
  Online. Association for Computational Linguistics.

\bibitem[{Chu et~al.(2017)Chu, Dabre, and Kurohashi}]{chu-etal-2017-empirical}
Chenhui Chu, Raj Dabre, and Sadao Kurohashi. 2017.
\newblock \href {https://doi.org/10.18653/v1/P17-2061} {An empirical comparison
  of domain adaptation methods for neural machine translation}.
\newblock In \emph{Proceedings of the 55th Annual Meeting of the Association
  for Computational Linguistics (Volume 2: Short Papers)}, pages 385--391,
  Vancouver, Canada. Association for Computational Linguistics.

\bibitem[{Dakwale and Monz(2017)}]{dakwale2017finetuning}
Praveen Dakwale and Christof Monz. 2017.
\newblock Finetuning for neural machine translation with limited degradation
  across in-and out-of-domain data.
\newblock \emph{Proceedings of the XVI Machine Translation Summit}, 117.

\bibitem[{Farajian et~al.(2017)Farajian, Turchi, Negri, and
  Federico}]{farajian-etal-2017-multi}
M.~Amin Farajian, Marco Turchi, Matteo Negri, and Marcello Federico. 2017.
\newblock \href {https://doi.org/10.18653/v1/W17-4713} {Multi-domain neural
  machine translation through unsupervised adaptation}.
\newblock In \emph{Proceedings of the Second Conference on Machine
  Translation}, pages 127--137, Copenhagen, Denmark. Association for
  Computational Linguistics.

\bibitem[{Freitag and Al-Onaizan(2016)}]{freitag2016fast}
Markus Freitag and Yaser Al-Onaizan. 2016.
\newblock Fast domain adaptation for neural machine translation.
\newblock \emph{arXiv preprint arXiv:1612.06897}.

\bibitem[{Gu et~al.(2017)Gu, Wang, Cho, and Li}]{DBLP:journals/corr/GuWCL17}
Jiatao Gu, Yong Wang, Kyunghyun Cho, and Victor O.~K. Li. 2017.
\newblock \href {http://arxiv.org/abs/1705.07267} {Search engine guided
  non-parametric neural machine translation}.
\newblock \emph{CoRR}, abs/1705.07267.

\bibitem[{Gu et~al.(2018)Gu, Wang, Cho, and Li}]{DBLP:conf/aaai/GuWCL18}
Jiatao Gu, Yong Wang, Kyunghyun Cho, and Victor O.~K. Li. 2018.
\newblock \href
  {https://www.aaai.org/ocs/index.php/AAAI/AAAI18/paper/view/17282} {Search
  engine guided neural machine translation}.
\newblock In \emph{Proceedings of the Thirty-Second {AAAI} Conference on
  Artificial Intelligence, (AAAI-18), the 30th innovative Applications of
  Artificial Intelligence (IAAI-18), and the 8th {AAAI} Symposium on
  Educational Advances in Artificial Intelligence (EAAI-18), New Orleans,
  Louisiana, USA, February 2-7, 2018}, pages 5133--5140. {AAAI} Press.

\bibitem[{Hasler et~al.(2021)Hasler, Domhan, Trenous, Tran, Byrne, and
  Hieber}]{hasler-etal-2021-improving}
Eva Hasler, Tobias Domhan, Jonay Trenous, Ke~Tran, Bill Byrne, and Felix
  Hieber. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.emnlp-main.666} {Improving
  the quality trade-off for neural machine translation multi-domain
  adaptation}.
\newblock In \emph{Proceedings of the 2021 Conference on Empirical Methods in
  Natural Language Processing}, pages 8470--8477, Online and Punta Cana,
  Dominican Republic. Association for Computational Linguistics.

\bibitem[{He et~al.(2021)He, Huang, Cui, Li, and Liu}]{he-etal-2021-fast}
Qiuxiang He, Guoping Huang, Qu~Cui, Li~Li, and Lemao Liu. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.acl-long.246} {Fast and
  accurate neural machine translation with translation memory}.
\newblock In \emph{Proceedings of the 59th Annual Meeting of the Association
  for Computational Linguistics and the 11th International Joint Conference on
  Natural Language Processing (Volume 1: Long Papers)}, pages 3170--3180,
  Online. Association for Computational Linguistics.

\bibitem[{Jalili~Sabet et~al.(2016)Jalili~Sabet, Negri, Turchi, C.~de Souza,
  and Federico}]{jalili-sabet-etal-2016-tmop}
Masoud Jalili~Sabet, Matteo Negri, Marco Turchi, Jos{\'e}~G. C.~de Souza, and
  Marcello Federico. 2016.
\newblock \href {https://doi.org/10.18653/v1/P16-4009} {{TM}op: a tool for
  unsupervised translation memory cleaning}.
\newblock In \emph{Proceedings of {ACL}-2016 System Demonstrations}, pages
  49--54, Berlin, Germany. Association for Computational Linguistics.

\bibitem[{Khandelwal et~al.(2021)Khandelwal, Fan, Jurafsky, Zettlemoyer, and
  Lewis}]{khandelwal2021nearest}
Urvashi Khandelwal, Angela Fan, Dan Jurafsky, Luke Zettlemoyer, and Mike Lewis.
  2021.
\newblock \href {https://openreview.net/forum?id=7wCBOfJ8hJM} {Nearest neighbor
  machine translation}.
\newblock In \emph{International Conference on Learning Representations}.

\bibitem[{Kingma and Ba(2015)}]{KingmaB14}
Diederik~P. Kingma and Jimmy Ba. 2015.
\newblock \href {http://arxiv.org/abs/1412.6980} {Adam: {A} method for
  stochastic optimization}.
\newblock In \emph{3rd International Conference on Learning Representations,
  {ICLR} 2015, San Diego, CA, USA, May 7-9, 2015, Conference Track
  Proceedings}.

\bibitem[{Kirkpatrick et~al.(2017)Kirkpatrick, Pascanu, Rabinowitz, Veness,
  Desjardins, Rusu, Milan, Quan, Ramalho, Grabska-Barwinska, Hassabis, Clopath,
  Kumaran, and Hadsell}]{kirkpatricEWC}
James Kirkpatrick, Razvan Pascanu, Neil Rabinowitz, Joel Veness, Guillaume
  Desjardins, Andrei~A. Rusu, Kieran Milan, John Quan, Tiago Ramalho, Agnieszka
  Grabska-Barwinska, Demis Hassabis, Claudia Clopath, Dharshan Kumaran, and
  Raia Hadsell. 2017.
\newblock \href {https://doi.org/10.1073/pnas.1611835114} {Overcoming
  catastrophic forgetting in neural networks}.
\newblock \emph{Proceedings of the National Academy of Sciences},
  114(13):3521--3526.

\bibitem[{Koehn and Knowles(2017)}]{koehn-knowles-2017-six}
Philipp Koehn and Rebecca Knowles. 2017.
\newblock \href {https://doi.org/10.18653/v1/W17-3204} {Six challenges for
  neural machine translation}.
\newblock In \emph{Proceedings of the First Workshop on Neural Machine
  Translation}, pages 28--39, Vancouver. Association for Computational
  Linguistics.

\bibitem[{Kudo and Richardson(2018)}]{DBLP:conf/emnlp/KudoR18}
Taku Kudo and John Richardson. 2018.
\newblock \href {https://doi.org/10.18653/v1/d18-2012} {Sentencepiece: {A}
  simple and language independent subword tokenizer and detokenizer for neural
  text processing}.
\newblock In \emph{Proceedings of the 2018 Conference on Empirical Methods in
  Natural Language Processing, {EMNLP} 2018: System Demonstrations, Brussels,
  Belgium, October 31 - November 4, 2018}, pages 66--71. Association for
  Computational Linguistics.

\bibitem[{Li et~al.(2022)Li, Su, Cai, Wang, and Liu}]{RATsurvey}
Huayang Li, Yixuan Su, Deng Cai, Yan Wang, and Lemao Liu. 2022.
\newblock \href {https://doi.org/10.48550/ARXIV.2202.01110} {A survey on
  retrieval-augmented text generation}.

\bibitem[{Luong and Manning(2015)}]{luong-manning-2015-stanford}
Minh-Thang Luong and Christopher Manning. 2015.
\newblock \href {https://aclanthology.org/2015.iwslt-evaluation.11} {{S}tanford
  neural machine translation systems for spoken language domains}.
\newblock In \emph{Proceedings of the 12th International Workshop on Spoken
  Language Translation: Evaluation Campaign}, pages 76--79, Da Nang, Vietnam.

\bibitem[{Meng et~al.(2021)Meng, Li, Zheng, Wu, Sun, Zhang, and
  Li}]{DBLP:journals/corr/abs-2105-14528}
Yuxian Meng, Xiaoya Li, Xiayu Zheng, Fei Wu, Xiaofei Sun, Tianwei Zhang, and
  Jiwei Li. 2021.
\newblock \href {http://arxiv.org/abs/2105.14528} {Fast nearest neighbor
  machine translation}.
\newblock \emph{CoRR}, abs/2105.14528.

\bibitem[{Neubig et~al.(2019)Neubig, Dou, Hu, Michel, Pruthi, and
  Wang}]{neubig-etal-2019-compare}
Graham Neubig, Zi-Yi Dou, Junjie Hu, Paul Michel, Danish Pruthi, and Xinyi
  Wang. 2019.
\newblock \href {https://doi.org/10.18653/v1/N19-4007} {compare-mt: A tool for
  holistic comparison of language generation systems}.
\newblock In \emph{Proceedings of the 2019 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics (Demonstrations)},
  pages 35--41, Minneapolis, Minnesota. Association for Computational
  Linguistics.

\bibitem[{Pham et~al.(2021)Pham, Crego, and Yvon}]{10.1162/tacl_a_00351}
MinhQuang Pham, Josep~Maria Crego, and François Yvon. 2021.
\newblock \href {https://doi.org/10.1162/tacl_a_00351} {{Revisiting
  Multi-Domain Machine Translation}}.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  9:17--35.

\bibitem[{Post(2018)}]{post-2018-call}
Matt Post. 2018.
\newblock \href {https://doi.org/10.18653/v1/W18-6319} {A call for clarity in
  reporting {BLEU} scores}.
\newblock In \emph{Proceedings of the Third Conference on Machine Translation:
  Research Papers}, pages 186--191, Brussels, Belgium. Association for
  Computational Linguistics.

\bibitem[{Robertson and Zaragoza(2009)}]{robertson-bm25-2009}
Stephen Robertson and Hugo Zaragoza. 2009.
\newblock \href {https://doi.org/10.1561/1500000019} {The probabilistic
  relevance framework: Bm25 and beyond}.
\newblock \emph{Found. Trends Inf. Retr.}, 3(4):333–389.

\bibitem[{Sachan et~al.(2021)Sachan, Patwary, Shoeybi, Kant, Ping, Hamilton,
  and Catanzaro}]{sachan-etal-2021-end}
Devendra Sachan, Mostofa Patwary, Mohammad Shoeybi, Neel Kant, Wei Ping,
  William~L. Hamilton, and Bryan Catanzaro. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.acl-long.519} {End-to-end
  training of neural retrievers for open-domain question answering}.
\newblock In \emph{Proceedings of the 59th Annual Meeting of the Association
  for Computational Linguistics and the 11th International Joint Conference on
  Natural Language Processing (Volume 1: Long Papers)}, pages 6648--6662,
  Online. Association for Computational Linguistics.

\bibitem[{Saunders et~al.(2019)Saunders, Stahlberg, de~Gispert, and
  Byrne}]{saunders-etal-2019-domain}
Danielle Saunders, Felix Stahlberg, Adri{\`a} de~Gispert, and Bill Byrne. 2019.
\newblock \href {https://doi.org/10.18653/v1/P19-1022} {Domain adaptive
  inference for neural machine translation}.
\newblock In \emph{Proceedings of the 57th Annual Meeting of the Association
  for Computational Linguistics}, pages 222--228, Florence, Italy. Association
  for Computational Linguistics.

\bibitem[{Swietojanski et~al.(2016)Swietojanski, Li, and
  Renals}]{swietojanski2016learning}
Pawel Swietojanski, Jinyu Li, and Steve Renals. 2016.
\newblock Learning hidden unit contributions for unsupervised acoustic model
  adaptation.
\newblock \emph{IEEE/ACM Transactions on Audio, Speech, and Language
  Processing}, 24(8):1450--1463.

\bibitem[{Thompson et~al.(2019)Thompson, Gwinnup, Khayrallah, Duh, and
  Koehn}]{thompson-etal-2019-overcoming}
Brian Thompson, Jeremy Gwinnup, Huda Khayrallah, Kevin Duh, and Philipp Koehn.
  2019.
\newblock \href {https://doi.org/10.18653/v1/N19-1209} {Overcoming catastrophic
  forgetting during domain adaptation of neural machine translation}.
\newblock In \emph{Proceedings of the 2019 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long and Short Papers)}, pages 2062--2068,
  Minneapolis, Minnesota. Association for Computational Linguistics.

\bibitem[{Vaswani et~al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones,
  Gomez, Kaiser, and Polosukhin}]{NIPS2017_7181}
Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones,
  Aidan~N Gomez, \L~ukasz Kaiser, and Illia Polosukhin. 2017.
\newblock \href
  {http://papers.nips.cc/paper/7181-attention-is-all-you-need.pdf} {Attention
  is all you need}.
\newblock In \emph{Advances in Neural Information Processing Systems 30}, pages
  5998--6008. Curran Associates, Inc.

\bibitem[{Vilar(2018)}]{vilar-2018-learning}
David Vilar. 2018.
\newblock \href {https://doi.org/10.18653/v1/N18-2080} {Learning hidden unit
  contribution for adapting neural machine translation models}.
\newblock In \emph{Proceedings of the 2018 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 2 (Short Papers)}, pages 500--505, New Orleans,
  Louisiana. Association for Computational Linguistics.

\bibitem[{Xia et~al.(2019)Xia, Huang, Liu, and Shi}]{Xia_Huang_Liu_Shi_2019}
Mengzhou Xia, Guoping Huang, Lemao Liu, and Shuming Shi. 2019.
\newblock \href {https://doi.org/10.1609/aaai.v33i01.33017297} {Graph based
  translation memory for neural machine translation}.
\newblock \emph{Proceedings of the AAAI Conference on Artificial Intelligence},
  33(01):7297--7304.

\bibitem[{Xu et~al.(2020)Xu, Crego, and Senellart}]{xu-etal-2020-boosting}
Jitao Xu, Josep Crego, and Jean Senellart. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.144} {Boosting
  neural machine translation with similar translations}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pages 1580--1590, Online. Association for
  Computational Linguistics.

\bibitem[{Zhang et~al.(2018)Zhang, Utiyama, Sumita, Neubig, and
  Nakamura}]{zhang-etal-2018-guiding}
Jingyi Zhang, Masao Utiyama, Eiichro Sumita, Graham Neubig, and Satoshi
  Nakamura. 2018.
\newblock \href {https://doi.org/10.18653/v1/N18-1120} {Guiding neural machine
  translation with retrieved translation pieces}.
\newblock In \emph{Proceedings of the 2018 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long Papers)}, pages 1325--1335, New Orleans,
  Louisiana. Association for Computational Linguistics.

\bibitem[{Zheng et~al.(2021)Zheng, Zhang, Guo, Huang, Chen, Luo, and
  Chen}]{zheng-etal-2021-adaptive}
Xin Zheng, Zhirui Zhang, Junliang Guo, Shujian Huang, Boxing Chen, Weihua Luo,
  and Jiajun Chen. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.acl-short.47} {Adaptive
  nearest neighbor machine translation}.
\newblock In \emph{Proceedings of the 59th Annual Meeting of the Association
  for Computational Linguistics and the 11th International Joint Conference on
  Natural Language Processing (Volume 2: Short Papers)}, pages 368--374,
  Online. Association for Computational Linguistics.

\end{thebibliography}

\pdfoutput=1

\documentclass[11pt]{article}

\usepackage{acl}
\usepackage{times}
\usepackage{latexsym}

\usepackage[T1]{fontenc}

\usepackage[utf8]{inputenc}

\usepackage{microtype}
\usepackage{times}
\usepackage{latexsym}
\usepackage{multirow}
\usepackage{url}

\usepackage{adjustbox}
\usepackage{graphicx}
\usepackage[scaled=.90]{helvet}
\usepackage{amsfonts}
\usepackage{float}
\restylefloat{table}
\usepackage{booktabs}
\usepackage{amsmath}
\usepackage{enumitem}
\usepackage{multirow}
\usepackage{subcaption}
\usepackage{cleveref}
\crefformat{section}{\S#2#1#3}
\crefformat{subsection}{\S#2#1#3}

\usepackage[normalem]{ulem}

\usepackage{caption}

\title{Improving Robustness of Retrieval Augmented Translation via \\ Shuffling of Suggestions}

\author{Cuong Hoang\thanks{\hspace{1.5mm}Work done while the authors were at AWS AI Labs.}, Devendra Sachan\footnotemark[1], Prashant Mathur, 
Brian Thompson, Marcello Federico\\ 
AWS AI Labs \\
\href{mailto:pramathu@amazon.com}{pramathu@amazon.com}}

\date{}

\begin{document}
\maketitle
\begin{abstract}
Several recent studies have reported dramatic performance improvements in neural machine translation (NMT) by augmenting translation at inference time with fuzzy-matches retrieved from a translation memory (TM).
However, these studies all operate under the assumption that the TMs available at test time are highly relevant to the testset. 
We demonstrate that for existing retrieval augmented translation methods, using a TM with a domain mismatch to the test set can result in substantially worse performance  compared to not using a TM at all. 
We propose a simple method to expose fuzzy-match NMT systems during training and show that it results in a system that is much more tolerant (regaining up to $5.8$ BLEU) to inference with TMs with domain mismatch. {Also, the model is still competitive to the baseline when fed with suggestions from relevant TMs.} %
\end{abstract}

\section{Introduction}
\label{sec:intro}
Retrieval Augmented Translation (RAT) refers to a paradigm 
which combines a translation model~\cite{NIPS2017_7181} with an external retriever module~\cite{RATsurvey}.
The \textit{retrieval module} (e.g. BM25 ranker \cite{robertson-bm25-2009} or a neural retriever \cite{cai-etal-2021-neural,sachan-etal-2021-end}) takes each source sentence as input and retrieves the top-$k$ most similar target translations from a Translation Memory (TM)~\cite{farajian-etal-2017-multi,DBLP:journals/corr/GuWCL17,bulte-tezcan-2019-neural}. 
The \textit{translation module} then encodes the input along with the top-$k$ fuzzy-matches, either by appending the suggestions to the input \cite{bulte-tezcan-2019-neural,xu-etal-2020-boosting} or using separate encoders for input and suggestions \cite{he-etal-2021-fast,cai-etal-2021-neural}. 
Decoder then either learns to copy or carry over the stylistics from the suggestions while generating the translation. 
In this work, we focus only on the translation module of this paradigm.

In existing literature, 
inference with RAT models have typically assumed that TMs are \textit{domain-matched} i.e. the test set is of the same domain as the translation memory. 
Many works (e.g. ~\citet{bulte-tezcan-2019-neural}, \citet{xu-etal-2020-boosting} and \citet{cai-etal-2021-neural}) have reported dramatic performance improvements under such setting. However, it is not clear how the models perform when there is a domain-mismatch between the TM and the test set.

In this work, we focus on the problem where the assumption of TM being domain-matched with test set does not hold. We explore the conditions where models are provided suggestions from a TM, but not from the same domain as the test set.
We show that RAT models suffer performance drop if fed with suggestions coming from less relevant TMs.

This finding is especially important from a usability standpoint. 
Often a translator may pick a TM as a best fit for a translation job if an ideal TM is not available or created just yet e.g. \texttt{IT} domain TM is picked for \texttt{Patent} translation job because the domains are closer and there no available TM for \texttt{Patent}. 
This could lead to multiple issues like ambiguous terminologies or multiple meanings across same context \cite{jalili-sabet-etal-2016-tmop,10.1007/s10590-016-9183-x}).
RAT model leveraging such (mismatched) TM ends up producing worse quality translation than a standard MT system.
Therefore, it is desirable that RAT models not only improve translation with suggestions from relevant TMs, but also be more robust to suggestions from less relevant TMs.

To this end, we propose an enhancement to the training of RAT models with a simple shuffling method to mitigate this problem. Instead of
always using $k$ most-relevant fuzzy-matches in training,
our method \textit{randomly samples} $k$ from a larger list 
(e.g. randomly sample 3 sentences from the top-$10$ matches). 
Our hypothesis is that if we systematically provide only top similar suggestions during training, the model will overly rely on the suggestions and simply copy the tokens in them. By shuffling the retrieved results, we ensure suggestions to be less similar to the input and train the system to be more robust to less relevant suggestions at test time. Our experiment results show that the model trained with shuffling of suggestions outperforms the standard RAT model by up to $+5.8$ BLEU score on average when suggestions come from less relevant TMs while dropping $-0.15$ BLEU on average when suggestions come from relevant TMs. 

To the best of our knowledge, this is the first work to consider the robustness of RAT methods, which we believe is critical for acceptance by human translators. 

\section{Related Work}
\label{sec:additional}

RAT is a form of domain adaptation, which is often achieved via continued training in NMT~\cite{freitag2016fast,luong-manning-2015-stanford}. 
However, RAT differs from standard domain adaptation techniques like continued training in that it is \textit{online}, that is, the model is not adapted during training and instead domain adaptation occurs at inference time. This makes RAT better suited for some real-world applications e.g. a single server with a single model loaded in memory can serve hundreds or thousands of users with custom translations, adapted to their unique TM. 
Other works have considered online adaptation outside the context of RAT, including \citet{vilar-2018-learning}, who propose Learning Hidden Unit Contributions  \cite{swietojanski2016learning} as a compact way to store many adaptations of the same general-domain model.

Previous works in retrieval augmented translation have mainly explored 
aspects of filtering fuzzy-matches by applying similarity thresholds  \cite{Xia_Huang_Liu_Shi_2019,xu-etal-2020-boosting}, leveraging word alignment information \cite{zhang-etal-2018-guiding,xu-etal-2020-boosting,he-etal-2021-fast} or re-ranking with additional score (e.g. word overlapping) \cite{DBLP:conf/aaai/GuWCL18,zhang-etal-2018-guiding}. 
Our approach do not make use of any filtering and as such do not require any ad-hoc optimization.
Our work is related to the use of $k$-nearest-neighbor for NMT  \cite{khandelwal2021nearest,zheng-etal-2021-adaptive} but it is less expensive and does not require storage and search over a large data store of context representations and corresponding target tokens \cite{DBLP:journals/corr/abs-2105-14528}.

Our work also relates to work in \textit{offline} adaptation which has addressed catastrophic forgetting 
of general domain knowledge during domain adaptation \cite{thompson-etal-2019-overcoming} via 
ensembling in-domain and out-of-domain models \cite{freitag2016fast}, 
mixing of in-domain and out-of-domain data during adaptation \cite{chu-etal-2017-empirical},
multi-objective learning and multi-output learning \cite{dakwale2017finetuning},
and 
elastic weight consolidation \cite{kirkpatricEWC, thompson-etal-2019-overcoming, saunders-etal-2019-domain}, 
or combinations of these techniques \cite{hasler-etal-2021-improving}.

\section{RAT with Shuffling}
\subsection{Retrieval Module}
We use Okapi BM25~\cite{robertson-bm25-2009}, a classical retrieval algorithm that performs search by computing lexical matches of the query with all sentences in the evidence, to obtain top-ranked sentences for each input.\footnote{To enable fast retrieval, we leverage the implementation provided by the ElasticSearch library, which can be found at \url{https://github.com/elastic/elasticsearch-py}.} Specifically, we built an index using source sentences of a TM. For every input, we collect top-$k$ (i.e. $k=\{1, 2, 3, 4, 5\}$  in our experiments) similar source side sentences and then use their corresponding target side sentences as fuzzy-matches.%

\subsection{Shuffling suggestions}\label{model_section}
We propose to relax the use of top-$k$ relevant suggestions during training by training RAT model with $k$ fuzzy-matches randomly sampled from a larger list. In our experiments, we sample $k$ from the top-$10$ matches, where top-$10$ is chosen based on our preliminary experiments. 

By shuffling retrieval fuzzy-matches, we ensure suggestions to be less similar to the target reference. With that, we expect the model learns to be more selective in using the suggestions for translation and thus be more robust to less relevant suggestions at test time. In fact, training models with noisy data has been shown to improve model's robustness to irrelevant data~\cite{DBLP:conf/iclr/BelinkovB18}.

\section{Data, Models \& Experiments}
\subsection{Data}
We conduct experiments in two language directions: En-De with five domain-specialized TMs\footnote{\texttt{Medical}, \texttt{Law}, \texttt{IT}, \texttt{Religion} and \texttt{Subtitles}.} and En-Fr with seven domain-specialized TMs.\footnote{\texttt{News}, \texttt{Medical}, \texttt{Bank}, \texttt{Law}, \texttt{IT}, \texttt{TED} and \texttt{Religion}.} En-De data is taken from \newcite{aharoni2020unsupervised}, which is a re-split version of the multi-domain data set from \newcite{koehn-knowles-2017-six} while En-Fr data set is taken from the multi-domain data set of \newcite{10.1162/tacl_a_00351}. We concatenate all parallel training data from multiple domains for training and similarly for development sets.

\subsection{Models and Training}
We experiment with a common RAT model used in \newcite{bulte-tezcan-2019-neural} and \newcite{xu-etal-2020-boosting}. The model encodes the input along with the top-$k$ fuzzy-matches by appending the suggestions to the input. We follow the recipe of \newcite{NIPS2017_7181} during training. We report translation quality with BLEU scores computed via \texttt{Sacrebleu}~\cite{post-2018-call}.\footnote{Details of all hyper-parameters of training and evaluation are presented in Appendix~\ref{appendix:params}.} We use \texttt{compare-mt} \cite{neubig-etal-2019-compare} for
significance testing with \texttt{bootstrap = 1000} and
\texttt{prob\_thresh = 0.05}. 

\subsection{Inference with TMs}
We run inference with RAT models using top-k suggestions coming from \textit{relevant} and \textit{less relevant} TMs. Specifically, for every input sentence in a test set from one specific domain, less relevant suggestions are retrieved from all TMs but not the TM coming from the same domain to the test set. Taking En-De and \texttt{IT} domain test set as an example, less relevant suggestions are collected from all TMs coming from \texttt{Medical}, \texttt{Law}, \texttt{Religion} and \texttt{Subtitles} but not \texttt{IT} domain.
\begin{table}[ht!]
\small
\centering
\begin{tabular}{l|l|c} 
\multicolumn{3}{c}{En $\rightarrow$ De}\\ \hline
Model &TM&\textbf{AVER}\\ \hline
Transformer &-&37.61\\\hline
RAT &\multirow{2}{*}{Less relevant}&32.89\\
RAT + Shuf.&&\textbf{35.28/+2.39}\\\hline
RAT &\multirow{2}{*}{Relevant}&41.34\\
RAT + Shuf.&&\textbf{41.19/-0.15}\\ \hline
\end{tabular}
\begin{tabular}{l|l|c}
\multicolumn{3}{c}{En $\rightarrow$ Fr}\\ \hline
Model &TM&\textbf{AVER}\\ \hline
Transformer&-&55.86\\ \hline
RAT &\multirow{2}{*}{Less relevant}&44.39\\
RAT + Shuf.&&\textbf{49.20/+5.81} \\ \hline
RAT &\multirow{2}{*}{Relevant}&56.59\\
RAT + Shuf.&&\textbf{56.44/-0.15}\\ \hline
\end{tabular}
\caption{Average BLEU scores of models across top-k values and across domains. The model trained with shuffling is more tolerant to domain mismatch in TMs while still being competitive when fed with suggestions from relevant TMs.}
\label{mainresults_average}
\end{table}
\begin{table*}[ht]
\small
\centering
\begin{tabular}{l|lllllll} 
\multicolumn{8}{c}{En $\rightarrow$ De}\\ \hline
Model & Top-k & IT & LAW & REL & MED & SUBT & \textbf{AVER}\\  \hline
Baseline  &-& 39.22&53.06&23.03&48.48&24.24&37.61\\  \hline
RAT &\multirow{2}{*}{k=2}&\textbf{41.16}&44.41&16.48&41.95&23.84&33.57\\
RAT + Shuf.&&40.81&\textbf{49.37*}&\textbf{21.01*}&\textbf{46.85*}&23.92&\textbf{36.39}\\\hline
RAT &\multirow{2}{*}{k=3}&\textbf{41.03}&43.61&15.11&40.35&23.35&32.69\\
RAT + Shuf.&&40.33&\textbf{46.52*}&\textbf{18.37*}&\textbf{44.66*}&\textbf{24.18}&\textbf{34.81}\\\hline
RAT &\multirow{2}{*}{k=4}&40.91&42.98&14.39&39.09&23.56&32.19\\
RAT + Shuf.&&\textbf{42.0}&\textbf{46.4*}&\textbf{18.23*}&\textbf{43.65*}&\textbf{24.02}&\textbf{34.86}\\\hline
\end{tabular}
\begin{tabular}{l|lllllllllll}
\multicolumn{10}{c}{En $\rightarrow$ Fr}\\ \hline
Model & Top-k& LAW & MED &IT & NEWS & BANK & REL & TED & \textbf{AVER}\\ \hline
Baseline &-& 65.58&57.92&48.89&34.43&57.32&86.87&40.03&55.86\\\hline
RAT & \multirow{2}{*}{k=2}&56.89&51.36&42.18&33.15&47.42&49.62&39.09&45.67\\
RAT + Shuf.&&\textbf{60.20*}&\textbf{52.66*}&\textbf{45.55*}&\textbf{34.12*}&\textbf{50.36*}&\textbf{81.76*}&\textbf{39.85}&\textbf{52.07}
\\\hline
RAT & \multirow{2}{*}{k=3}&56.48&50.31&40.41&33.25&45.32&43.61&38.31&43.96\\
RAT + Shuf.&& \textbf{57.81*}&\textbf{51.89*}&\textbf{42.78*}&\textbf{33.51}&\textbf{48.08*}&\textbf{73.11*}&\textbf{38.75}&\textbf{49.42}\\\hline
RAT & \multirow{2}{*}{k=4}&54.93&49.98&39.77&\textbf{33.05}&45.03&39.26&37.75&42.82\\
RAT + Shuf. && \textbf{56.33*}&\textbf{50.92*}&\textbf{42.07*}&32.94&\textbf{46.58}&\textbf{60.31*}&\textbf{38.48}&\textbf{46.80}\\
\hline
\end{tabular}
\caption{Model comparison with \textbf{less relevant} suggestions. The best BLEU for models with a specific top-k value is highlighted in bold, and "*" indicates whether the models are statistical significance difference with the others.}\label{mainresults_less relevant}
\end{table*}
\begin{table*}[ht]
\small
\centering
\begin{tabular}{l|lllllll} 
\multicolumn{8}{c}{En $\rightarrow$ De}\\ \hline
Model & Top-k & IT & LAW & REL & MED & SUBT & \textbf{AVER}\\  \hline
Baseline  &-& 39.22&53.06&23.03&48.48&24.24&37.61\\ \hline
RAT & \multirow{2}{*}{k=2}&\textbf{41.52}&\textbf{57.86}&30.67&\textbf{53.52}&23.83&\textbf{41.48}\\
RAT + Shuf.&&40.78&57.82&\textbf{30.87}&53.34&\textbf{24.07}&41.38\\\hline
RAT &\multirow{2}{*}{k=3}&\textbf{41.43}&\textbf{57.67}&\textbf{31.24}&\textbf{53.05*}&23.66&\textbf{41.41}\\
RAT + Shuf.&&40.90&56.98&30.65&51.36&\textbf{24.24}&40.83\\\hline
RAT &\multirow{2}{*}{k=4}&41.35&57.54&31.7&52.43&23.53&41.31\\
RAT + Shuf.&&\textbf{42.24}&\textbf{58.90*}&\textbf{33.02*}&\textbf{53.7}&\textbf{23.94}&\textbf{42.36}\\\hline
\end{tabular}
\begin{tabular}{l|lllllllllll}
\multicolumn{10}{c}{En $\rightarrow$ Fr}\\ \hline
Model & Top-k& LAW & MED &IT & NEWS & BANK & REL & TED & \textbf{AVER}\\ \hline
Baseline &-& 65.58&57.92&48.89&34.43&57.32&86.87&40.03&55.86\\\hline
RAT& \multirow{2}{*}{k=2}&\textbf{70.25*}&\textbf{60.20*}&\textbf{50.28}&33.95&\textbf{58.65*}&83.99&40.32&\textbf{56.81}\\
RAT + Shuf.&&69.45&59.42&49.41&\textbf{34.38}&57.46&\textbf{84.15}&\textbf{40.63}&56.41
\\\hline
RAT&\multirow{2}{*}{k=3}&\textbf{70.70}&\textbf{61.02}&\textbf{52.57}&33.61&\textbf{60.67}&84.84&39.47&57.55\\
RAT + Shuf.&&{70.03}&{60.41}&51.69&\textbf{33.64}&60.41&\textbf{85.15}&\textbf{39.66}& \textbf{57.28}
\\\hline
RAT &\multirow{2}{*}{k=4}&\textbf{70.07}&\textbf{60.23*}&\textbf{50.28}&33.20&\textbf{57.5}&84.65&39.37&\textbf{56.47}\\
RAT + Shuf.&& 69.73&59.03&49.4&\textbf{33.32}&56.63&\textbf{85.29}&\textbf{39.78}&56.17\\\hline
\end{tabular}
\caption{Model comparison with \textbf{relevant} suggestions. The
best BLEU for models with a specific top-k value is highlighted in bold, and "*" indicates whether the models are statistical significance difference with the others.}\label{mainresults_relevant}
\end{table*}
\section{Results}\label{multidomain}
\subsection{Main results}
We experiment with comprehensive top-k values of $\{1, 2, 3, 4, 5\}$ and compare models when fed with suggestions from both relevant and less relevant TMs. The summary results across top-k values and across domains are collected in Table~\ref{mainresults_average} for both En-De and En-Fr.
When the standard RAT models are fed with less relevant suggestions, we see a drop of 12.5\% (37.61~$\rightarrow$~32.89) for En-De and 20.5\% (55.86~$\rightarrow$~44.39) for En-Fr against standard Transformer based MT system. This shows how brittle the RAT models are when the assumption -  TMs are of same domain as test set - breaks.  
Our approach with shuffling closes the gap by 50\% from 4.72 (37.61~$\rightarrow$~32.89) to 2.39 (37.61~$\rightarrow$~35.28) for En-De and 54.7\% similarly for En-Fr when RAT models are fed with less relevant suggestions while dropping 0.15 when the suggestions are from relevant TMs. This is an acceptable trade-off from the point of view of a user because of the significant gains observed in the less relevant TM setting.

Details with a subset of top-$k$ values of $\{2, 3, 4\}$ are in Table~\ref{mainresults_less relevant} (with less relevant TMs) and Table~\ref{mainresults_relevant} (with relevant TMs).\footnote{We report a subset of results due to space constraint. Comprehensive results are shown in Appendix~\ref{appendix:results}.}
The improvements with shuffling are consistent for all $k$s when models are fed less relevant TMs and clearly support that shuffling suggestions during training significantly improves robustness of RAT models.
In Table~\ref{mainresults_relevant}, where the models are fed with relevant TMs, we observe a drop in translation quality in some cases (e.g. \texttt{MED} domain in En-De with k=3) but on an average performance of the models across domains remains competitive. %

\begin{table}[ht!]
\small
\begin{tabular}{l|llllllllllll} 
Model &IT&LAW&REL&MED&SUBT\\  \cline{0-8}
\multicolumn{6}{c}{Less Relevant TM}\\ \cline{0-5}
RAT &24.82&17.81&19.25&14.73&20.69\\
+ Shuf.&23.51*&17.04*&17.74*&13.42*&20.05*\\\hline
\multicolumn{6}{c}{Relevant TM}\\ \cline{0-5}
RAT &27.01&40.75&36.37&39.45&20.67&\\
+ Shuf.&25.63*&39.05*&34.33*&37.46*&19.99*&\\\hline
\end{tabular}
\caption{Average percentage of suggestion tokens that appear in MT outputs produced by RAT models across five En-De test sets, with both relevant and less relevant TMs. ``*'' indicates statistical significance difference.}\label{tab:percentage}
\end{table}

\subsection{Usage of Suggestions}
To understand the direct impact of suggestions on translation, we looked at the percentage of tokens in suggestions that appear in output produced by model. Average percentage of overlapping tokens for each test set is in Table~\ref{tab:percentage}, with $k$ = 3 and En-De as a case study. 

Our results show a consistently significantly less use of less relevant TMs by RAT models trained with shuffling method. This clearly support our postulation that RAT models trained with shuffling are more selective in using suggestions for translation, which explains the robustness of the model when it is run inference with less relevant suggestions.

As the model is more selective, it also tends to less use of suggestions from relevant TMs as in Table~\ref{tab:percentage}. This explains the translation accuracy drop in some cases (e.g. \texttt{MED} domain in En-De with k=3) when fed with relevant suggestions in exchange of the robustness.

\section{Conclusion}
We show that standard RAT systems are not robust to domain-mismatch at test time. We proposed a training procedure of shuffling fuzzy-match suggestions to mitigate the problem. Our experiment results show that the model trained with shuffling significantly outperforms the one without shuffling, across multiple domains on two language pairs when suggestions come from less relevant TMs at inference time. We also show that RAT with shuffling is still competitive to models without shuffling across all domains when fed with relevant suggestions. Further analysis shows that the model trained with shuffling is more selective in using suggestions for translation, which explains the increased robustness.

\bibliography{main.bbl}
\bibliographystyle{acl_natbib}

\clearpage

\addcontentsline{toc}{section}{Appendices}
\renewcommand{\thesubsection}{\Alph{subsection}}
\section*{Appendix}

\subsection{Model Parameters}\label{appendix:params}
We employed Transformer big architecture with 6 encoder and 6 decoder layers. Hidden size was set to  $1024$ and maximum input length to $1024$ tokens. We employed a joint source-target language subword vocabulary of size $32K$ using Sentencepiece algorithm~\cite{DBLP:conf/emnlp/KudoR18}.

We use the Adam optimizer~\cite{KingmaB14} with $\beta_1 = 0.9$, $\beta_2 = 0.98$ and $\epsilon = 10^{-9}$; and (ii) increase the learning rate linearly for the first $4K$ training steps and decrease it thereafter; (iii) use batch size of $32K$ source tokens and $32K$ target tokens. Checkpoints are saved after every $10K$ iterations during training. We use dropout of $0.1$ and label-smoothing of $0.1$. We train models with maximum of $300K$ iterations.

For evaluation we use Sacrebleu with hash value of \texttt{nrefs:1|case:mixed|eff:no|tok:13a|\\smooth:exp|version:2.0.0}.

\subsection{Full results with $k=\{1, 2, 3, 4, 5\}$}\label{appendix:results}
Details with full set of top-$k$ values of $\{1, 2, 3, 4, 5\}$ are in Table~\ref{amainresults_less relevant} (i.e. with less relevant TMs) and Table~\ref{amainresults_relevant} (i.e. with relevant TMs).

\begin{table*}[ht]
\small
\begin{tabular}{l|llllllllllll} 
\multicolumn{9}{c}{En $\rightarrow$ De}\\ \cline{0-8}
Model &TM&Top-k&IT&LAW&REL&MED&SUBT&\textbf{AVER}\\  \cline{0-8}
Baseline  &-&-& 39.22&53.06&23.03&48.48&24.24&37.61\\  \cline{0-8}
RAT &\multirow{10}{*}{Less relevant}&\multirow{2}{*}{k=1}&\textbf{40.82}&45.61&17.98&42.65&24.08&34.22\\
RAT + Shuf.&&&39.56&\textbf{50.89*}&\textbf{21.6*}&\textbf{47.8*}&\textbf{24.33}&\textbf{36.84}\\\cline{3-9}
RAT &&\multirow{2}{*}{k=2}&\textbf{41.16}&44.41&16.48&41.95&23.84&33.57\\
RAT + Shuf.&&&40.81&\textbf{49.37*}&\textbf{21.01*}&\textbf{46.85*}&23.92&\textbf{36.39}\\\cline{3-9}
RAT &&\multirow{2}{*}{k=3}&\textbf{41.03}&43.61&15.11&40.35&23.35&32.69\\
RAT + Shuf.&&&40.33&\textbf{46.52*}&\textbf{18.37*}&\textbf{44.66*}&\textbf{24.18}&\textbf{34.81}\\\cline{3-9}
RAT &&\multirow{2}{*}{k=4}&40.91&42.98&14.39&39.09&23.56&32.19\\
RAT + Shuf.&&&\textbf{42.0}&\textbf{46.4*}&\textbf{18.23*}&\textbf{43.65*}&\textbf{24.02}&\textbf{34.86}\\\cline{3-9}
RAT &&\multirow{2}{*}{k=5}&40.83&41.93&13.8&38.55&23.67&31.76\\
RAT + Shuf.&&&\textbf{41.04}&\textbf{44.41*}&\textbf{16.25*}&\textbf{41.57*}&\textbf{24.12}&\textbf{33.48}\\\hline
\multicolumn{9}{c}{En $\rightarrow$ Fr}\\ \hline
Model &TM&Top-k& LAW & MED &IT & NEWS & BANK & REL & TED & \textbf{AVER}\\ \hline
Baseline &-&-& 65.58&57.92&48.89&34.43&57.32&86.87&40.03&55.86\\\hline
RAT&\multirow{10}{*}{Less relevant}&\multirow{2}{*}{k=1}&58.3&51.54&44.99&33.51&48.49&59.62&39.07&47.93\\
RAT + Shuf.&&&\textbf{60.97*}&\textbf{53.48*}&\textbf{45.95}&\textbf{33.76}&\textbf{51.2}&\textbf{82.37*}&\textbf{39.73}&\textbf{52.49}\\\cline{3-11}
RAT&&\multirow{2}{*}{k=2}&56.89&51.36&42.18&33.15&47.42&49.62&39.09&45.67\\
RAT + Shuf.&&&\textbf{60.20*}&\textbf{52.66*}&\textbf{45.55*}&\textbf{34.12*}&\textbf{50.36*}&\textbf{81.76*}&\textbf{39.85}&\textbf{52.07}
\\\cline{3-11}
RAT&&\multirow{2}{*}{k=3}&56.48&50.31&40.41&33.25&45.32&43.61&38.31&43.96\\
RAT + Shuf.&&&\textbf{57.81*}&\textbf{51.89*}&\textbf{42.78*}&\textbf{33.51}&\textbf{48.08*}&\textbf{73.11*}&\textbf{38.75}&\textbf{49.42}\\\cline{3-11}
RAT&&\multirow{2}{*}{k=4}&54.93&49.98&39.77&\textbf{33.05}&45.03&39.26&37.75&42.82\\
RAT + Shuf.&&&\textbf{56.33*}&\textbf{50.92*}&\textbf{42.07*}&32.94&\textbf{46.58}&\textbf{60.31*}&\textbf{38.48}&\textbf{46.80}
\\\cline{3-11}
RAT&&\multirow{2}{*}{k=5}&54.36&49.91&39&32.2&44.38&33.88&37.23&41.57\\
RAT + Shuf.&&&\textbf{55.43*}&\textbf{50.41}&\textbf{40.51*}&\textbf{32.81}&\textbf{45.26}&\textbf{54.02*}&\textbf{37.91}&\textbf{45.19}\\\cline{3-11}
\end{tabular}
\caption{Full results of comparing models with \textit{less relevant} fuzzy-matches. Here, top-k values are with $\{1, 2, 3, 4, 5\}$. The
best BLEU for models with a specific top-k value is highlighted in bold, and "*" indicates whether the models are statistical significance difference with the others.}\label{amainresults_less relevant}
\end{table*}
\begin{table*}[ht]
\small
\begin{tabular}{l|llllllllllll} 
\multicolumn{9}{c}{En $\rightarrow$ De}\\ \cline{0-8}
Model &TM&Top-k&IT&LAW&REL&MED&SUBT&\textbf{AVER}\\  \cline{0-8}
Baseline  &-&-& 39.22&53.06&23.03&48.48&24.24&37.61\\  \cline{0-8}
RAT &\multirow{10}{*}{Relevant}&\multirow{2}{*}{k=1}&\textbf{41.24*}&\textbf{57.66*}&\textbf{29.11*}&\textbf{53.73*}&\textbf{24.04}&\textbf{41.16}\\
RAT + Shuf.&&&39.55&55.84&25.92&50.8&24.30&39.28\\\cline{3-9}
RAT &&\multirow{2}{*}{k=2}&\textbf{41.52}&\textbf{57.86}&30.67&\textbf{53.52}&23.83&\textbf{41.48}\\
RAT + Shuf.&&&40.78&57.82&\textbf{30.87}&53.34&\textbf{24.07}&41.38\\\cline{3-9}
RAT &&\multirow{2}{*}{k=3}&\textbf{41.43}&\textbf{57.67}&\textbf{31.24}&\textbf{53.05*}&23.66&\textbf{41.41}\\
RAT + Shuf.&&&40.90&56.98&30.65&51.36&\textbf{24.24}&40.83\\\cline{3-9}
RAT &&\multirow{2}{*}{k=4}&41.35&57.54&31.7&52.43&23.53&41.31\\
RAT + Shuf.&&&\textbf{42.24}&\textbf{58.90*}&\textbf{33.02*}&\textbf{53.7}&\textbf{23.94}&\textbf{42.36}\\\cline{3-9}
RAT &&\multirow{2}{*}{k=5}&40.58&57.2&32.1&53.21&23.69&41.36\\
RAT + Shuf.&&&\textbf{41.43}&\textbf{58.32*}&\textbf{33.07}&\textbf{53.89}&\textbf{23.78}&\textbf{42.10}\\\hline
\multicolumn{9}{c}{En $\rightarrow$ Fr}\\ \hline
Model &TM&Top-k& LAW & MED &IT & NEWS & BANK & REL & TED & \textbf{AVER}\\ \hline
Baseline &-&-& 65.58&57.92&48.89&34.43&57.32&86.87&40.03&55.86\\\hline
RAT&\multirow{10}{*}{Relevant}&\multirow{2}{*}{k=1}&\textbf{70.39*}&\textbf{59.45}&\textbf{51.77*}&34.07&\textbf{57.42}&81.15&39.93&56.31\\
RAT + Shuf.&&&69.2&58.83&49.75&\textbf{34.26}&57.29&\textbf{84.92*}&\textbf{40.16}&\textbf{56.34}
\\\cline{3-11}
RAT&&\multirow{2}{*}{k=2}&\textbf{70.25*}&\textbf{60.20*}&\textbf{50.28}&33.95&\textbf{58.65*}&83.99&40.32&\textbf{56.81}\\
RAT + Shuf.&&&69.45&59.42&49.41&\textbf{34.38}&57.46&\textbf{84.15}&\textbf{40.63}&56.41
\\\cline{3-11}
RAT&&\multirow{2}{*}{k=3}&\textbf{70.70}&\textbf{61.02}&\textbf{52.57}&33.61&\textbf{60.67}&84.84&39.47&57.55\\
RAT + Shuf.&&&{70.03}&{60.41}&51.69&\textbf{33.64}&60.41&\textbf{85.15}&\textbf{39.66}& \textbf{57.28}
\\\cline{3-11}
RAT&&\multirow{2}{*}{k=4}&\textbf{70.07}&\textbf{60.23*}&\textbf{50.28}&33.20&\textbf{57.5}&84.65&39.37&\textbf{56.47}\\
RAT + Shuf.&&&69.73&59.03&49.4&\textbf{33.32}&56.63&\textbf{85.29}&\textbf{39.78}&56.17\\\cline{3-11}
RAT&&\multirow{2}{*}{k=5}&68.99&59.39&\textbf{49.61}&32.85&\textbf{56.92}&83.88&39.14&55.83\\
RAT + Shuf.&&&\textbf{69.55}&\textbf{59.51}&49.60&\textbf{33.37}&56.45&\textbf{84.13}&\textbf{39.44}&\textbf{56.01}
\\\cline{3-11}
\end{tabular}
\caption{Full results of comparing models with \textit{relevant} fuzzy-matches. Here, top-k values are with $\{1, 2, 3, 4, 5\}$. The
best BLEU for models with a specific top-k value is highlighted in bold, and "*" indicates whether the models are statistical significance difference with the others.}\label{amainresults_relevant}
\end{table*}

\end{document}